OpenStudy (anonymous):

I really need help with this question please help me. The Smith family wants to order a pizza. A large cheese pizza cost $10. There is an additional charge of $1.50 per topping. Which equation below represents this function. a) f(x)= 10+1.50x b) f(x)= 11.50x c) f(x)= 1.50+10x

OpenStudy (anonymous):

ur equation a and c are the same i guess.

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

the fixed cost is $10.... our you will always pay that plus $1.50per topping so let x be the number of toppings the cost f(x) = 1.5x + 10 e.g 2 topping f(2) = 1.5*2 + 10 =13